OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 9x2 − 6x + 1 36b2 − 24b + 8 16x2 + 24x + 9 4a2 − 10a + 25

OpenStudy (anonymous):

all but 2nd?

OpenStudy (anonymous):

does 9x2 mean \[9x ^{2}\] ?

OpenStudy (anonymous):

9x2-6x+1=(3x-1)^2 36b^2-24b-8= i dont even want to do it 16x2+24x+9=(4x+3)^2 4a2-10a+25=(2a-5)^2

OpenStudy (anonymous):

and yes it does usually